OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (5, 9); Slope: 2 A. y – 5 = 2(x + 9) B. y + 9 = 2(x – 5) C. y – 9 = 2(x + 5) D. y – 9 = 2(x – 5)

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) so when given Point: (5, 9); Slope: 2 plug in m = 2 x1 = 5 and y1 = 9 into y - y1 = m(x - x1)

OpenStudy (anonymous):

d.

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thx I got 5 more questions though

jigglypuff314 (jigglypuff314):

I can help with those too :)

OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–8, 3); Slope: 6 A. y + 3 = 6x – 48 B. y + 3 = 6(x + 8) C. y – 3 = 6(x – 8) D. y – 3 = 6(x + 8)

jigglypuff314 (jigglypuff314):

same idea as before :) point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) so when given Point: (–8, 3); Slope: 6 plug in m = 6 x1 = -8 and y1 = 3 into y - y1 = m(x - x1)

OpenStudy (anonymous):

c.

jigglypuff314 (jigglypuff314):

be careful :) when you plug in the x1 = -8 y - y1 = m(x - (-8)) ^ negative times negative equals positive

OpenStudy (anonymous):

ohhh b.

jigglypuff314 (jigglypuff314):

the y1 = 3 would remain negative so you would have gotten y - 3 = 6(x + 8)

OpenStudy (anonymous):

d. definitely d.

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thx :D

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

2 more instead of 3

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: \[-\frac{ 2 }{3} \]

jigglypuff314 (jigglypuff314):

same idea as before :) point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) so when given Point: (1, –7); Slope: −2/3 plug in m = -2/3 x1 = 1 and y1 = -7 into y - y1 = m(x - x1)

OpenStudy (anonymous):

\[y-7= -\frac{2}{3}\]

OpenStudy (anonymous):

\[y+7=\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

\[y-7=\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

\[y+7=-\frac{ 2 }{ 3 }\]

jigglypuff314 (jigglypuff314):

\[y - (-7) = (-\frac{ 2 }{ 3 }) (x - (1))\]

OpenStudy (anonymous):

d.

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

:D one last question

jigglypuff314 (jigglypuff314):

ok :)

OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks:

OpenStudy (anonymous):

Which equation gives the cost, y, in terms of the number of pages, x? A. y = 83.3x + 0.65 B. y = 0.012x + 49.985 C. y = 0.012x + 0.65 D. y = 0.012x + 1.25

jigglypuff314 (jigglypuff314):

1.85 minus 1.25 = 0.6 so every 50 pages = 0.6 then 0.6 divided by 50 = how much one page costs = 0.012 when pages = 0 cost = 1.25 - 0.6 = 0.65 so when y = mx + b m = 0.012 and b = 0.65

OpenStudy (anonymous):

c

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thx! :D

jigglypuff314 (jigglypuff314):

you're welcome :)